OpenStudy (anonymous):

MEDAL FOR THE FIRST TO ANSWER Which expressions could complete this equation so that it has an infinite number of solutions? 4(3x - 5) + 6 = ____ Choose exactly 2 answers A.4x - 14 + 8x B. 12x - 14 C.6x + 2(3x - 8) D.7(x - 2)+6x

OpenStudy (mayankdevnani):

what do you think ?

OpenStudy (anonymous):

This one i dont know?

OpenStudy (mayankdevnani):

lolzz

OpenStudy (mayankdevnani):

lets put option B first, \[\large \bf 12x-20+6=12x-14\] \[\large \bf 12x-12x=-14+20-6=0=0=\color{red}{Infinite~no.of~solutions}\]

OpenStudy (mayankdevnani):

just like this...solve your question and pick up another answer :)

OpenStudy (mayankdevnani):

understood ? @Kyliejones

OpenStudy (anonymous):

B. is the closest Real answer: 12x -11

OpenStudy (mayankdevnani):

lolzz...you are wrong !!! @sop002

OpenStudy (mayankdevnani):

just see my solution and correct your answer :)

OpenStudy (anonymous):

Back sorry

OpenStudy (anonymous):

@mayankdevnani follow order of operation

OpenStudy (anonymous):

B. and C.

OpenStudy (mayankdevnani):

see :- if \[\large \bf 0=constant----->No~solution\]

OpenStudy (mayankdevnani):

so @Kyliejones understood ??

OpenStudy (anonymous):

Guys im running out of time whats the answer?

OpenStudy (mayankdevnani):

can you show your work??? @sop002 to prove your answer

OpenStudy (mayankdevnani):

option B is one of the answer

OpenStudy (mayankdevnani):

and option C is other answer :)

OpenStudy (mayankdevnani):

@Kyliejones

OpenStudy (anonymous):

Thanks so much!!

OpenStudy (mayankdevnani):

welcome :)

OpenStudy (anonymous):

correct !!!!!

OpenStudy (mayankdevnani):

thnx... ^^^

OpenStudy (triciaal):

A and B

OpenStudy (anonymous):

I realized that my answered was wrong. #sorry @mayankdevnani

OpenStudy (mayankdevnani):

its ok....you are studying here,,,.... but i am happy that you realized your mistake :)

OpenStudy (mayankdevnani):

and REMEMBER :- \[\large \bf 0=0---->Infinite~no.of~solutions\] \[\large \bf 0=constant--->No~Solution\]

OpenStudy (mayankdevnani):

@sop002

OpenStudy (mayankdevnani):

and also WELCOME TO OPEN STUDY :)

OpenStudy (anonymous):

: )

OpenStudy (anonymous):

she got it rong

OpenStudy (mayankdevnani):

no way !!!!

OpenStudy (anonymous):

she told me

OpenStudy (mayankdevnani):

oh!!!! option A is correct :)

OpenStudy (mayankdevnani):

sorry...everyone !!!

OpenStudy (anonymous):

yes

OpenStudy (mayankdevnani):

EXPLANATION :- \[\large \bf 12x-20+6=4x-14+8x\] \[\large \bf 0=0=\color{red}{Infinite~no.of~Solutions}\]

OpenStudy (mayankdevnani):

sorry :(... @Kyliejones and @sop002

OpenStudy (triciaal):

better approach and correct 4(3x - 5) + 6 = ____ 12 x - 20 + 6 = 12 x -14 to have infinite number of solutions x is not a fixed number find which options are equivalent to 12 x - 14 to start option B simplify the others option A also give 12 x - 14 so equivalent are options A and B

OpenStudy (mayankdevnani):

and thanks @lsherron

OpenStudy (mayankdevnani):

yeah... i have corrected already... @triciaal

OpenStudy (anonymous):

your welcome